OpenStudy (anonymous):

Simplify the given expression to rational exponent form and justify each step by identifying the properties of rational exponents used. All work must be shown. 1 over the cube root of the quantity of x to the negative sixth power

OpenStudy (anonymous):

\[\frac{1}{\sqrt[3]{x^{-6}}}\]?

OpenStudy (anonymous):

since \(\frac{-6}{3}=-2\) you can start with \[\frac{1}{x^{-2}}\] then take the reciprocal

OpenStudy (anonymous):

ok how do i show all the work

OpenStudy (anonymous):

on paper

OpenStudy (anonymous):

ok thanks ill take a picture of it

OpenStudy (anonymous):

can you help me with another?

OpenStudy (anonymous):

Simplify the given expression to radical form and justify each step by identifying the properties of rational exponents used. All work must be shown. x to the two–thirds power, over x to the four–ninths power